OpenStudy (anonymous):

A triangle has the following measurements. What is a possible length for the third side? AB = 4, BC = 17, CA = ?

OpenStudy (anonymous):

is bc = 17 the longest side?

OpenStudy (anonymous):

AC < 4 + 17 AC<21 AC > 17-4 AC>13

OpenStudy (anonymous):

CA = AB^2 + BC^2 = 4^2 + 17^2 = 305 \[\sqrt{305}\] = 17,46